OpenStudy (anonymous):

simplify the expression: cot^2x/cscx+1

OpenStudy (lgbasallote):

this seems familiar...well are you familiar that \[\large cot^2{x} = csc^2{x} - 1\] so let's sub that \[\large \frac {csc^2{x} - 1}{cscx + 1}\] \[\large \frac{(cscx+1)(cscx-1)}{cscx+1}\] i think you can see the rest now? :DD

OpenStudy (lgbasallote):

if you have any questions regarding what i did..please feel free to ask :)

OpenStudy (anonymous):

i got it ^^ thx